OpenStudy (anonymous):

Can you help me with this one? What is the formula for sinA+sinB in Sum of two functions ?

OpenStudy (anonymous):

in(A+B)=sin A cos B + cos A sin B sin(A-B)=sin A cos B - cos A sin B cos(A+B)=cos A cos B - sin A sin B cos(A-B)=cos A cos B + sin A sin B

OpenStudy (unklerhaukus):

\[\large\text{Sum & Difference Formulas}\] \[\sin(u ± v) = \sin (u) \cos (v) ± \cos (u) \sin (v)\] \[\cos(u ± v) = \cos u \cos v ∓ \sin u \sin v\] \[\tan(u±v)= \frac{\tan u±\tan v }{1∓\tan u\tan v}\]

OpenStudy (anonymous):

wow!!! Thank you for replying!!!!!! :D

OpenStudy (anonymous):

k